OpenStudy (anonymous):

solve the following systems of equations by substitution. y= x^2+1 y= 4x+1

OpenStudy (anonymous):

Solve\[ x^2 +1 = 4 x +1\\ x^2 =4x\\ x(x-4) =0 \\x=0,\quad x=4 \] Replace in either equations to find y